# Dual-Domain Homogeneous Fusion with Cross-Modal Mamba and Progressive Decoder for 3D Object Detection

### v1

@@ IV-B Results and Comparison
Method Voxel Stride mAP NDS
(1) [1] 69.6 71.9
@@ IV-C Ablation Studies
Ablation of each module As shown in Table V, we evaluate the contribution of each module to the detection performance. We replace the decoder in BEVFusion with deformable attention as the baseline. First, the proposed voxel fusion network significantly improves performance by 1.6 mAP and 0.9 NDS. Introducing only the BEV fusion network also provides a gain of 0.7 mAP and 0.5 NDS. As reported in the third row, the combination of both homogeneous fusion networks achieves a substantial improvement of 2.2 mAP and 1.4 NDS. Additionally, the progressive query generation module enhances object detection recall, contributing 0.7 mAP. Finally, integrating voxel decoder into the original BEV decoder to form the progressive decoder further increases performance by 0.3 mAP and 0.7 NDS. In summary, the four proposed components fully exploit the potential of dual-domain multi-modal features for high-performance object detection.
@@
TABLE X: Comparsion of image sizes and channels of LiDAR backbone.
Model Image Size or Channels mAP NDS
Image Backbone 256$\times$704
384$\times$1056
LiDAR Backbone (16,32,64,128)
(32,64,128,256)
@@
Analysis of homogeneous voxel fusion network As shown in Table IV, we analyze the role of the Mamba-based modules in the voxel fusion stage. In the second row, we introduce a 3D U-Net with only IV-Mamba, without directly fusing image and LiDAR voxels. This approach improves performance by 1.2 mAP and 0.5 NDS, demonstrating that IV-Mamba can effectively capture long-range dependencies within each modality. The CV-Mamba can also enhance network performance, indicating the effectiveness of global fusion of voxels with different modalities and distributions. By combining both approaches with a simple merge-and-split strategy, we not only align multi-modal features but also preserve the relative independence of each modality, avoiding interference from noise or low-quality information from other modality, thereby further improving accuracy. As shown in Table V, setting the voxel strides to [1, 2, 4] in HVF achieves the best performance, as increasing the downsampling rate allows the network to capture structural information at different granularities.
@@
Analysis of homogeneous BEV fusion network We further investigate HBF. As shown in Table VI, the SS2D module in the IB-Mamba enables full-scene feature propagation across the dense BEV map in four directions, resulting in improvements of 0.3 mAP and 0.1 NDS. Our proposed CB-Mamba achieves gains of 0.5 mAP and 0.2 NDS. It effectively propagates multi-modal features while adaptively exploring their latent relationships, thereby eliminating feature misalignment in the BEV domain. Notably, the combination of IB-Mamba and CB-Mamba yields the best performance, which benefits subsequent the query generation and decoder process.
@@
Comparsion of different activation method In PQG, we explore different methods to activate BEV features for extracting harder queries in the second stage. As shown in Table VII, following the approach described in [48], utilizing a residual block yields a minor improvement of 0.2 mAP and 0.1 NDS. Our proposed HIA module further incorporates spatial cues from easy queries, enabling the identification of remaining hard objects in driving scenarios and achieving higher accuracy gains. However, it is important to note that the cross-attention between queries and BEV features introduces additional computational overhead, which limits the number of queries and stages. We plan to address this issue in the future work.
@@
Analysis of progressive decoder In the progressive decoder, both BEV-based and voxel-based decoders are employed for query refinement. As shown in Table IX, we explore various configurations of their layer counts (M and N) to analyze their impact on performance. In the first row, using only a single voxel-based decoder yields suboptimal accuracy, even underperforming the second row, which adopts a 3-layer BEV decoder as in [48]. The BEV map at the end of the network contains richer contextual information and higher-level semantic features compared to voxels. Additionally, its dense representation is more robust to point cloud density variations than sparse voxel representation, both of which are critical for object detection in complex environments. However, as shown in the third row, adding a voxel decoder after the BEV decoder improves performance by 0.2 mAP and 0.7 NDS, indicating that incorporating voxel features with fine geometric details can enhance regression accuracy of predicted boxes. However, further increasing layer counts provides negligible performance gains. In Table X, we demonstrate that the mixing block and self-attention mechanism can adaptively aggregate shape features and long-range spatial information respectively, each contributing to performance improvements. As reported in the last row, combining them achieves the best results.
Analysis of on different scaling choices
@@ IV-D Comparsion of Different Homogeneous Fusion Methods
First, we evaluate their performance under different weather and lighting conditions. Since DeepInteraction++ and UVTR do not provide relevant data in their papers, we reproduce their models using open-source code to obtain the accuracy. As shown in Table XI, DDHFusion achieves the highest mAP across all environments. Unlike UVTR, DDHFusion is not negatively affected by rainy conditions and performs better even in poor lighting environments. This robustness stems from the comprehensive intra-modal and cross-modal feature learning in DDHFusion, which progressively filters out corrupt information to learn robust latent representations. Additionally, as shown in Table XII, we evaluate the performance of different models on object detection at varying distances, where “Near,” “Middle,” and “Far” represent ranges of $<20$m, [20[20[ 20m, $30$m], and $>30$m, respectively. Again, our proposed DDHFusion outperforms other methods across all distance ranges.
@@
Figure 8 showcases the detection results of DDHFusion, demonstrating its ability to accurately classify and localize objects of various categories in complex environments. Figure 9 provides a qualitative comparison with other models. As shown in the first row, DDHFusion successfully detects three distant and crowded objects. In the second row, bounding boxes generated by DDHFusion show minimal deviations. In the third row, it successfully identifies the small and distant object that other methods fail to detect. These results highlight the superiority of our dual-domain homogeneous fusion approach.
@@
Our method demonstrates competitive performance in terms of both memory consumption and runtime, highlighting its efficiency and practicality. For a fair comparison, all models are deployed in the mmdetection framework and tested on a single 4090D GPU. As shown in Table XIII, The computational resource consumption of DDHFusion is affordable, demonstrating the efficiency of our Mamba-based fusion modules. GraphBEV is the most lightweight, as it performs fusion only in the BEV domain. In contrast, UVTR incurs the highest computational burden due to its retention of dense voxels, while DDHFusion mitigates this issue through voxel sampling operations. Notably, DeepInteraction++ employs flash-attention for acceleration. Similarly, we believe that further hardware optimization of Mamba could significantly improve efficiency.

### v2

@@ IV-B Results and Comparison
As shown in Table I, we report the results of DDHFusion on the NuScenes test set. It achieves an impressive 73.8 mAP and 75.6 NDS without model ensemble or test-time augmentation, surpassing all other state-of-the-art methods. Compared to the second-best Bico-Fusion model, DDHFusion shows improvements of 1.4 mAP and 1.1 NDS. In Table II, we also present the performance of DDHFusion on the NuScenes validation set. DDHFusion outperforms the second-best DeepInteraction++ model by 1.0 mAP and 0.5 NDS. By increasing parameters and expanding the input image size, DDHFusion-Large achieves even greater accuracy advantages. These improvements are attributed to the complementary strengths voxel and BEV homogeneous domains, as well as advanced mechanism that we propose to provide larger receptive fields, fewer false negatives and higher localization accuracy.
@@
Method Downsampling strides mAP NDS
(1) [1] 69.5 71.9
@@ IV-C Ablation Studies
Ablation of each module. As shown in Table V, we evaluate the contribution of each module to the detection performance. We replace the decoder in BEVFusion with deformable attention as the baseline. First, the proposed HVF significantly improves performance by 1.6 mAP and 1.1 NDS. Introducing only the HBF network also provides a gain of 0.7 mAP and 0.6 NDS. As reported in the third row, the combination of both homogeneous fusion networks achieves a substantial improvement of 2.4 mAP and 1.6 NDS. Additionally, the PQG module enhances detection recall, contributing 0.6 mAP and 0.4 NDS. Finally, integrating voxel decoder into the original BEV decoder to form the progressive decoder further increases performance by 0.2 mAP and 0.7 NDS. In summary, the four proposed components can fully exploit the potential of the dual-domain fusion framework for high-performance object detection.
@@
TABLE X: Analysis of scaling choices.
Scaling choice mAP NDS
Baseline 71.6 73.8
+ Doubling channels in LiDAR backbone 71.9 74.0
+ Expanding input image size 72.8 74.4
@@
Analysis of homogeneous voxel fusion network. As shown in Table IV, we analyze the role of the Mamba-based modules in the voxel fusion stage. In the second row, we introduce a 3D U-Net with only IV-Mamba, without directly fusing image and LiDAR voxels. It achieves a performance improvement of 1.0 mAP and 0.7 NDS, which can be attributed to IV-Mamba’s capability to effectively capture long-range dependencies within each modality. The CV-Mamba can also enhance network performance, indicating the effectiveness of global fusion of voxel features with different modalities and distributions. By combining both IV-Mamba and CV-Mamba with the merge-and-split operation, we not only align multi-modal features but also preserve the relative independence of each modality, avoiding interference from noise or low-quality information from other modality, thereby further improving accuracy. As shown in Table V, setting the downsampling strides to $[1,2,4]$ in HVF can achieve the best performance, as increasing the downsampling rate allows the network to learn structural information at different granularities.
@@
Analysis of homogeneous BEV fusion network We further investigate HBF. As shown in Table VI, the SS2D module in IB-Mamba enables global feature diffusion across the dense BEV map in four directions, resulting in improvements of 0.3 mAP and 0.1 NDS. Our proposed CB-Mamba achieves gains of 0.5 mAP and 0.2 NDS. It effectively propagates multi-modal features while adaptively exploring their latent relationships and eliminating feature misalignment in the BEV domain. As reported in the last row, the combination of IB-Mamba and CB-Mamba yields the best performance, which benefits the subsequent query generation and feature decoding.
@@
Comparsion of different activation method In PQG, we choose different methods to activate BEV features for extracting harder queries in the second stage. As reported in Table VII, utilizing a residual block yields a minor improvement. Our proposed HIA module further incorporates spatial cues from easy queries to facilitate the identification of remaining hard objects in driving scenarios and achieve higher accuracy gains. However, it is important to note that the cross-attention between queries and BEV features introduces additional computational overhead, which limits the number of queries and stages. We plan to address this issue in future work.
@@
Analysis of progressive decoder In the progressive decoder, both BEV-based and voxel-based decoders are employed for query refinement. As shown in Table IX, we explore various configurations of their layer counts (M and N) to analyze their impact on performance. In the first row, using only a single voxel-based decoder yields suboptimal accuracy, even underperforming the second row, which adopts a three-layer BEV decoder. This is because the BEV map at the end of the network contains richer contextual information and higher-level semantic features compared to voxels. Additionally, the BEV dense representation is more robust to point density variations than sparse voxel representation. Therefore, BEV features are critical for object detection in complex environments. However, as shown in the third row, adding a voxel decoder after the BEV decoder further improves performance by 0.2 mAP and 0.7 NDS, indicating that incorporating voxel features with fine geometric details can enhance regression accuracy of predicted boxes. However, further increasing M provides negligible performance gains. In Table IX, we demonstrate that the mixing block and self-attention can adaptively aggregate 3D shape features and long-range spatial information respectively, each contributing to performance improvements. As reported in the last row, combining them can achieve the best result.
Analysis of on different scaling choices Following previous state-of-the-art methods, we explore increasing the model parameters or enlarging the input image size to achieve better accuracy. As indicated in Table X, doubling the channels in the LiDAR backbone only provides a marginal improvement of 0.3 mAP and 0.2 NDS. In contrast, expanding the input image size from $256\times 704$ to $384\times 1056$ yields a boost of 0.9 mAP and 0.4 NDS. This demonstrates that providing higher-quality image voxels $V_{I}$ and BEV features $B_{I}$ is effective in enhancing the network’s performance.
@@ IV-D Comparsion of Different Homogeneous Fusion Methods
First, we evaluate their performance under different weather and lighting conditions. Since DeepInteraction++ and UVTR do not provide relevant data in their papers, we reproduce their models using open-source codes to obtain their accuracy. As shown in Table XI, DDHFusion achieves the highest mAP across all environments. Unlike UVTR, DDHFusion is not negatively affected by rainy conditions and performs better even in poor lighting environments. This robustness stems from the comprehensive intra-modal and cross-modal feature learning in DDHFusion, which progressively filters out corrupt information to learn robust latent representations. Additionally, as shown in Table XII, we evaluate the performance of different models on object detection at varying distances, where “Near,” “Middle,” and “Far” represent ranges of $<20m$, $[20m,30m]$, and $>30m$, respectively. Again, our proposed DDHFusion outperforms other methods across all distance ranges.
@@
Fig. 8 showcases the detection results of DDHFusion, demonstrating its ability to accurately classify and localize objects of various categories in complex environments. Fig. 9 provides a qualitative comparison with other models. As shown in the first row, DDHFusion successfully detects three distant and crowded objects, demonstrating its ability to learn more discriminative representations from the two domains. In the second row, the bounding boxes generated by DDHFusion exhibit minimal deviations, which can be attributed to the rich structural details derived from the voxel domain. In the third row, it successfully identifies a small and distant object that other methods fail to detect, proving the effectiveness of the PQG module. These results highlight the superiority of our dual-domain homogeneous fusion framework.
As reported in Table XIII, we also compare their memory consumption and inference speeds. To ensure a fair comparison, all models are implemented within the mmdetection3D framework and tested on a single 4090D GPU. As shown in the last row, the computational resource consumption of DDHFusion is acceptable, which can be attributed to the efficiency of our Mamba-based fusion modules. GraphBEV is the most lightweight, as it performs fusion only in the BEV domain. In contrast, UVTR incurs the highest computational burden due to its retention of dense voxels, while DDHFusion mitigates this issue through the SAFS module. It is worth noting that DeepInteraction++ utilizes flash-attention for acceleration. Similarly, we believe that further hardware optimization of Mamba could lead to significant efficiency improvements.
@@ V Conclusion
In this work, we propose DDHFusion, which innovatively combines voxel and BEV homogeneous fusion paradigms and achieves exceptional performance. Another significant contribution is the successful extension of the recently popular Mamba architecture to the multi-modal 3D detection task, fully leveraging its versatility. Specifically, we employ cross-modal Mamba structures based on voxel or BEV representations in the HVF and HBF networks, which significantly enhance the receptive field during the fusion process and address the issue of feature misalignment. In the feature decoding stage, we also introduce PQG and PD to improve the target recall and regression accuracy of the network. In the future, we plan to explore the corresponding temporal fusion pipeline to further unleash the potential of DDHFusion.